OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. -2x+5/3y=-5 3x=7/2y=9/2 Show work PLEASE :) i know y is -3 can you help me find x?

OpenStudy (anonymous):

y = -3 -2x +(5/3)(-3) = -5 -2x - 15/3 = -5 -2x - 5 = -5 -2x = 0 x = 0

OpenStudy (anonymous):

Thanks :)